Two in-phase loudspeakers, which emit sound in all directions, are sitting side by side. One of them is moved sideways by 2.0 m, then forward by 8.0 m. Afterward, constructive interference is observed 1/4 1/2, and 3/4 the distance between the speakers along the line that joins them, and at no other positions along this line.

 What is the maximum possible wavelength of the sound waves?
Answer #1

We know,

λ = 2 L

√ [(2)2 + (8)2] = √(4 + 64) = √68 = 8.25 m

The distance between the spot is:

L = 3/4 * (8.25) - 1/4 * (5)

= 6.188 - 1.25

= 4.94 m

The longest wavelength between the spots is,

λ = 2 L

= 2 * 4.94

= 9.88 m